Assume that you know that the aggregate production function exhibits constant returns to scale, that there are only two factors, capital and income, and that

  • aggregate capital income = $1.25 billion
  • aggregate labour income = $4 billion
  • capital stock growth rate = 1.5 percent
  • employment growth rate = 2.3 percent
  • output growth rate = 2.7 percent

then the rate of growth of total factor productivity equals

Select one:

a. -0.9 percent.

b. 0.6 percent.

c. 2.7 percent.

d. there is insufficient information with which to answer this question.
Answer #1

We are given that aggregate capital income = $1.25 billion and aggregate labour income = $4 billion

Hence share of capital in income = 1.25 / (4 + 1.25) = 0.238 and share of labor in income = 4 / (1.25 + 4) = 0.762

Growth rate of output = share of capital in income * capital stock growth rate + share of labor in income * employment growth rate + growth rate of TFP

0.027 = 0.238*0.015 + 0.762*0.023 + growth rate of TFP

growth rate of TFP = 0.59% OR 0.6%

Select B